# XLIP: Cross-modal Attention Masked Modelling for Medical Language-Image Pre-Training

### v2

@@ Abstract
Vision-and-language pretraining (VLP) in the medical field utilizes contrastive learning on image-text pairs to achieve effective transfer across tasks. Yet, current VLP approaches with the masked modelling strategy face two challenges when applied to the medical domain. First, current models struggle to accurately reconstruct key pathological features due to the scarcity of medical data. Second, most methods only adopt either paired image-text or image-only data, failing to exploit the combination of both paired and unpaired data. To this end, this paper proposes a XLIP (Masked modelling for medical Language-Image Pre-training) framework to enhance pathological learning and feature learning via unpaired data. First, we introduce the attention-masked image modelling (AttMIM) and entity-driven masked language modelling module (EntMLM), which learns to reconstruct pathological visual and textual tokens via multi-modal feature interaction, thus improving medical-enhanced features. The AttMIM module masks a portion of the image features that are highly responsive to textual features. This allows XLIP to improve the reconstruction of highly similar image data in medicine efficiency. Second, our XLIP capitalizes unpaired data to enhance multimodal learning by introducing disease-kind prompts. The experimental results show that XLIP achieves SOTA for zero-shot and fine-tuning classification performance on five datasets. Our code will be available at https://github.com/White65534/XLIP.
{IEEEkeywords}
@@
Medical Language-Image Pretraining, Mask Image Modeling, attention, zero-shot
1 Introduction
@@
Vision-and-language pretraining (VLP) has garnered increasing attention in the medical field, primarily due to its capability to transfer representations effectively across various downstream tasks, including zero- or few-shot recognition. This ability significantly reduces reliance on extensive annotated data and efficiently detects various pathologies. Most VLP approaches emphasize contrastive learning to understand the relationship between visual and linguistic elements [30] Masked modelling has progressed and yielded promising results in pure visual [10, 4, 44], and textual [18] self-supervised representation learning. Masked image modelling (MIM) reconstructs masked image segments using surrounding context, while masked language modelling (MLM) predicts masked words in sentences through adjacent context, both enhancing comprehension in their respective visual and linguistic domains. The success has inspired some researchers to explore the significance of MIM and MLM within the scope of VLP. This involves randomly masking elements in a single modality [24, 34] or across multiple modalities [20, 6] during VLP, as shown in Figure 1 A.
@@
Figure 1: Modules A and B illustrate the distinctions between the existing VLP training frameworks and our proposed approach. Methods in module A only use the random mask strategy, and the reconstruction process does not involve multi-modal interaction. Module B shows that our method takes advantage of multi-modal interaction and uses an attention mechanism to guide the mask.
@@
Applying MIM and MLM to the medical domain still presents significant challenges, primarily due to two key limitations. First, traditional random masking misaligns with the specific needs of medical data, which fails to capture pathological representations critical for clinical diagnosis. In MIM, random masking inadequately captures the subtle but critical variations in anatomical structures, crucial for precise medical diagnosis. Concurrently, in MLM, this approach can lead to a diminished focus on accurately predicting medical entities, generating less medically informative content. The latest research has started to explore attention-guided MIM [37, 17], which are prone to finding crucial image regions for masking instead of random masking. However, constrained by the scarcity of medical data, models still struggle to accurately identify important medical areas during masking. Second, MIM and MLM were initially conceived for single-modality self-supervised learning, not accounting for the strong interplay between medical images and clinical texts. This oversight limits their effectiveness in VLP scenarios where cross-modal interactions are vital. The advancement [42, 8] has proven that integrating modality interaction in MIM/MLM can enhance the recognition ability of meaningful regions, but heavily depending on the medical image-text pairs.
@@
This motivates us to propose novel cross-modal attention Masked modelling for medical Language-Image Pre-training, called XLIP. Our XLIP is designed to accomplish three self-supervised learning tasks, as shown in Figure 1 B. First, we design a novel attention-masked image modelling module (AttMIM) for medical images. It detects discriminative features to mask by harnessing feature interactions between vision and language modalities. Specifically, our AttMIM blends masks generated through image self-attention, image-report cross-attention, and prompt-driven attention. While the first two modules mine instance-level interactions, the prompt-driven attention exploits global-level interactions between images and common disease terms, independent of paired data. As such, even having missing corresponding reports, our prompt-driven attention still localizes pathological features by leveraging the global affinity with frequent diseases. Secondly, we propose the entity-driven masked language modelling module (EntMLM) for medical reports, which masks informative entities and reconstructs the masked words. We also incorporate the learned image representations into the text decoder to aid the model in understanding medical entities with greater detail and nuance. Besides, we employ standard contrastive learning to align medical images and reports. To solve the inconsistent convergence rate problem between contrastive learning and masked modelling objectives, we first warm up the masked modelling tasks for specific iterations and then jointly train three tasks end-to-end.
@@
Our XLIP is pre-trained on two large-scale medical datasets, one is MIMIC-CXR with paired chest X-ray images and reports, and the other is the PadChest with only chest X-ray images. The learned image representations are transferred to five downstream classification datasets under zero-shot and fine-tuning settings, all of which achieved SOTA respectively.
@@
2 Related Works
2.1 Medical Vision Language Pretraining
Driven by the effectiveness of self-supervised pre-training methods in NLP and CV, there is promising potential for customizing VLP methods to medical imaging analysis, such as lesion detection [26] and segmentation [40, 51, 50]. The most recent research exploring VLP is now centred around: improving feature extraction capabilities [27, 22]; improving model structure [23, 33]; improving training methods [22, 7]; improving model compatibility with different modalities [43]; introducing preprocessing modules [28]; introducing prior knowledge [41]. As an application and extension of VLP in the medical field, Medical VLP focuses on understanding the content of medical images and texts. The latest outstanding research mainly focuses on the following three directions: 1) Model structures, particularly improvements in dual encoders [15, 45] and fusion encoders [22, 19, 35, 46]; 2) Scaling up training data, by collecting high-quality medical multi-modal datasets from various platforms to enhance representational learning [12, 25, 2]; 3) utilization of medical-specific prior knowledge to enhance model representational performance through prompts [41, 39].
2.2 Mask Image/Language Modelling
The development of MIM in computer vision, represented by MAE [10], SimMIM [44], and iBOT [53], marks a significant advance in self-supervised learning. MAE and SimMIM innovate in image masking, while iBOT combines these advances with the pre-training technology of self-distillation. Recent approaches have extensively explored various fusions of SSL techniques. For example, CTITM [6] utilizes MIM and MLM to augment CLIP, thereby enhancing the expressive power of the visual coder. MaskVLM [20] not only effectively integrates CLIP, MIM, and MLM, but also innovatively explores strategies for images and text to guide each other in generating masks, thus optimizing the synergy between modalities. However, these methods have limitations: they either don’t make full use of the available multimodal information, depending instead on features from a single mode to perform tasks, or their exploration of masking strategies is confined to using random methods.
2.3 Attention Masked modelling
@@
In the realm of attention-based mask modeling, hard patch mining (HPM) stands out as a groundbreaking training paradigm for MIM, designed to enhance a model’s comprehension of image content [37]. HPM method first has the model generate a challenging mask and then trains the model to predict the masked patches, similar to traditional methods. This approach enables the model to learn where it is worth applying masks while simultaneously learning how to solve the problems. Also, HPM introduces new metrics to measure the difficulty of the pre-training task by reconstructing the loss, which could help the model to focus more on challenging problems.
Meanwhile, traditional random masking strategies, effective in general domains, fall short in medical contexts where precision is crucial [8, 11, 29, 48, 41, 52]. MedIM [42] mitigates the unique challenges posed by healthcare data by improving mask strategies.
However, current methods either rely solely on single-mode features for their attention mechanisms [20] or have limited interaction between multimodal features, such as only employing feature matching [42]. Our proposed method solves these problems well by employing cross-attention for multimodal feature interaction while performing attention-masked modelling on features that incorporate multimodal information, which allows the model to understand image and text information in a more integrated way.
Figure 2: Our XLIP framework builds upon CLIP, integrating MIM and MLM modules, with a redesigned masking strategy to enhance model representation. The key contributions include three simple yet effective designs: generating masks through feature interaction, fusing masks to augment adaptability, and refining the text masking strategy with Medical Entity Recognition. Incorporating these designs into our multimodal pre-training framework significantly boosts the model’s zero-shot performance.
3 Methodology
Figure 2 illustrates the pipeline of the proposed XLIP. XLIP excels in representation learning by using image-report contrastive learning and masked modelling across both paired and unpaired medical image-report data. Central to XLIP are two innovative modules: Attention-Masked Image Modeling (AttMIM) and Entity-Driven Masked Language Modeling (EntMLM).
@@ 3 Methodology
EntMLM, on the other hand, concentrates on masking and reconstructing significant entities within medical reports, utilizing insights from image representations. This integration facilitates a more nuanced and detailed comprehension of medical terminology, enhancing the capabilities of both visual and textual learners within the XLIP framework.
3.1 Image and Text Encoders
3.1.1 Image encoder
We employ the Vision Transformer (ViT) [9] with a patch size of 16 as the image encoder, initialized with the weights of the clip-base, to process the given input medical images $x$. Initially, $x$ is transformed into a sequence of flattened patches. These patches are subsequently embedded and introduced into a 12-layer transformer followed by a linear projection layer, denoted as $F(\cdot)$ and $Linear(\cdot)$, resulting in encoded representations of visual tokens:
$E_{\text{img}}=Linear(F(x))\in\mathbb{R}^{N_{\text{img}}\times C},$ (1)
@@ 3.1.2 Text encoder
$E_{\text{report}}=Linear(T(t_{report}))\in\mathbb{R}^{N_{\text{report}}\times C},$ (2)
$E_{\text{prompt}}=Linear(T(t_{prompt}))\in\mathbb{R}^{N_{\text{prompt}}\times C},$ (3)
@@
3.2 Attention-masked Image Modeling Module
@@
3.2.1 Attentions extraction
Figure 3: Different masking strategies result in varying concerns regarding mask application. Masks that are specifically tailored to the pathological characteristics of the lesion area are more effective than those applied randomly.
@@ 3.2.1 Attentions extraction
$A_{r}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{report}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{report}},$ (4)
@@
$A_{p}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{prompt}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{prompt}},$ (5)
@@
Image self-attention directs the model to focus on the most challenging parts of the image features during convergence, which tend to be more difficult regions [37], thus improving the model’s ability and efficiency in understanding the image [49, 14]. Based on the self-attention mechanism, the image features after interacting with the image features themselves can be obtained as follows:
$A_{i}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{img}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{img}},$ (6)
@@
3.2.2 Attention-based mask generation and blending.
We introduce the function $M(A,\lambda)$, which retrieves indices corresponding to the top values in feature matrix $A$, with the number of indices determined as a proportion $\lambda_{1}$ of the matrix’s size. This function effectively transforms these indexes into a mask corresponding to the input matrix, which is the high-activated attention masking results in the attention extraction. For more details, please refer to Fig in the appendix. We blend the masks produced by different strategies in a union manner and define this result as $M_{b}$, formulated as:
$M_{b}=M(Ar;\lambda_{1})\cup M(Ap;\lambda_{1})\cup M(Ai;\lambda_{1}).$ (7)
@@
$\mathbb{M}_{\text{AttMIM}}={\rm RandomSubSet}(M_{b},\lambda_{2}).$ (8)
@@
${\rm MIM}_{\text{input}}=\mathbb{M}_{\text{AttMIM}}\cdot E_{\text{img}}.$ (9)
@@
3.3 Entity-driven Masked Language Modeling Module

### v3

# MMCLIP: Cross-modal Attention Masked Modelling for Medical Language-Image Pre-Training

@@ Abstract.
Vision-and-language pretraining (VLP) in the medical field utilizes contrastive learning on image-text pairs to achieve effective transfer across tasks. To further enhance visual and textual representation learning, many recent approaches adopt masked modeling strategies that randomly hide input tokens during training. Nevertheless, the adoption of masked modeling strategies in existing VLP methods gives rise to two key challenges in medical applications. First, current models struggle to accurately reconstruct key pathological features due to the scarcity of medical data. Second, most methods only adopt either paired image-text or image-only data, failing to exploit the combination of both paired and unpaired data. To this end, this paper proposes the MMCLIP (Masked Medical Contrastive Language-Image Pre-Training) framework to enhance pathological learning and feature learning via unpaired data. First, we introduce the attention-masked image modelling (AttMIM) and entity-driven masked language modelling module (EntMLM), which learns to reconstruct pathological visual and textual tokens via multi-modal feature interaction, thus improving medical-enhanced features. The AttMIM module masks a portion of the image features that are highly responsive to textual features. This allows MMCLIP to improve the reconstruction of highly similar image data in medicine efficiency. The EntMLM module identifies and masks key medical entities in the text with Named Entity Recognition (NER), and reconstructs them with support from visual features, enabling richer understanding of disease-related language. Second, our MMCLIP capitalizes unpaired data to enhance multimodal learning by introducing disease-kind prompts. The experimental results show that MMCLIP achieves SOTA for zero-shot and fine-tuning classification performance on five datasets. Our code will be available at https://github.com/White65534/MMCLIP.
@@
Medical Language-Image Pretraining, Mask Image Modeling, Self-supervised Learning
ccs: Computing methodologies Image representations
1. Introduction
@@
Vision-and-language pretraining (VLP) has garnered increasing attention in the medical field, primarily due to its capability to transfer representations effectively across various downstream tasks, including zero- or few-shot recognition. This ability significantly reduces reliance on extensive annotated data and efficiently detects various pathologies. Most VLP approaches emphasize contrastive learning to understand the relationship between visual and linguistic elements (of Radiology et al., 2014) Masked modelling has progressed and yielded promising results in pure visual (He et al., 2022; Bao et al., 2021; Xie et al., 2022a), and textual (Kenton and Toutanova, 2019) self-supervised representation learning. Masked image modelling (MIM) reconstructs masked image segments using surrounding context, while masked language modelling (MLM) predicts masked words in sentences through adjacent context, both enhancing comprehension in their respective visual and linguistic domains. The success has inspired some researchers to explore the significance of MIM and MLM within the scope of VLP. This involves randomly masking elements in a single modality (Li et al., 2023; Sun et al., 2023) or across multiple modalities (Kwon et al., 2022; Chen et al., 2023) during VLP, as shown in Figure 1 A.
@@
Figure 1. Modules A and B illustrate the distinctions between the existing VLP training frameworks and our proposed approach. Methods in module A only use the random mask strategy, and the reconstruction process does not involve multi-modal interaction. Module B shows that our method takes advantage of multi-modal interaction and uses an attention mechanism to guide the mask.
@@
Applying MIM and MLM to the medical domain still presents significant challenges, primarily due to two key limitations. First, traditional random masking misaligns with the specific needs of medical data, which fails to capture pathological representations critical for clinical diagnosis. In MIM, random masking inadequately captures the subtle but critical variations in anatomical structures, crucial for precise medical diagnosis. Concurrently, in MLM, this approach can lead to a diminished focus on accurately predicting medical entities, generating less medically informative content. The latest research has started to explore attention-guided MIM (Wang et al., 2023; Kakogeorgiou et al., 2022), which are prone to finding crucial image regions for masking instead of random masking. However, constrained by the scarcity of medical data, models still struggle to accurately identify important medical areas during masking. Second, MIM and MLM were initially conceived for single-modality self-supervised learning, not accounting for the strong interplay between medical images and clinical texts. This oversight limits their effectiveness in VLP scenarios where cross-modal interactions are vital. The advancement (Xie et al., 2023; Chen et al., 2022) has proven that integrating modality interaction in MIM/MLM can enhance the recognition ability of meaningful regions, but heavily depending on the medical image-text pairs.
@@
This motivates us to propose novel multimodal Masked Medical Contrastive Language-Image Pre-Training, named MMCLIP. Our MMCLIP is designed to accomplish three self-supervised learning tasks, as shown in Figure 1 B. First, we design a novel attention-masked image modelling module (AttMIM) for medical images. It detects discriminative features to mask by harnessing feature interactions between vision and language modalities. Specifically, our AttMIM blends masks generated through image self-attention, image-report cross-attention, and prompt-driven attention. While the first two modules mine instance-level interactions, the prompt-driven attention exploits global-level interactions between images and common disease terms, independent of paired data. As such, even having missing corresponding reports, our prompt-driven attention still localizes pathological features by leveraging the global affinity with frequent diseases. Secondly, we propose the entity-driven masked language modelling module (EntMLM) for medical reports, which masks informative entities and reconstructs the masked words. We also incorporate the learned image representations into the text decoder to aid the model in understanding medical entities with greater detail and nuance. Besides, we employ standard contrastive learning to align medical images and reports. To solve the inconsistent convergence rate problem between contrastive learning and masked modelling objectives, we first warm up the masked modelling tasks for specific iterations and then jointly train three tasks end-to-end.
@@
Our MMCLIP is pre-trained on two large-scale medical datasets, one is MIMIC-CXR with paired chest X-ray images and reports, and the other is the PadChest with only chest X-ray images. The learned image representations are transferred to five downstream classification datasets under zero-shot and fine-tuning settings, all of which achieved SOTA respectively.
@@
2. Related Works
2.1. Medical Vision Language Pretraining
Driven by the effectiveness of self-supervised pre-training methods in NLP and CV, there is promising potential for customizing VLP methods to medical imaging analysis, such as lesion detection (Liu et al., 2019) and segmentation (Wu et al., 2023a; Zhang et al., 2023b, a). The most recent research exploring VLP is now centred around: improving feature extraction capabilities (Lu et al., 2019; Li et al., 2019); improving model structure (Li et al., 2020; Su et al., 2019); improving training methods (Li et al., 2019; Chen et al., 2020); improving model compatibility with different modalities (Xie et al., 2022b); introducing preprocessing modules (Mu et al., 2022); introducing prior knowledge (Wu et al., 2023b). As an application and extension of VLP in the medical field, Medical VLP focuses on understanding the content of medical images and texts. The latest outstanding research mainly focuses on the following three directions: 1) Model structures, particularly improvements in dual encoders (Jia et al., 2021; Yao et al., 2021) and fusion encoders (Li et al., 2019; Kim et al., 2021; Tan and Bansal, 2019; Yu et al., 2021); 2) Scaling up training data, by collecting high-quality medical multi-modal datasets from various platforms to enhance representational learning (Huang et al., 2023; Lin et al., 2023; Awais et al., 2023); 3) utilization of medical-specific prior knowledge to enhance model representational performance through prompts (Wu et al., 2023b; Wang et al., 2022).
2.2. Mask Image or Language Modelling
The development of MIM in computer vision, represented by MAE (He et al., 2022), SimMIM (Xie et al., 2022a), and iBOT (Zhou et al., 2021), marks a significant advance in self-supervised learning. MAE and SimMIM innovate in image masking, while iBOT combines these advances with the pre-training technology of self-distillation. Recent approaches have extensively explored various fusions of SSL techniques. For example, CTITM (Chen et al., 2023) utilizes MIM and MLM to augment CLIP, thereby enhancing the expressive power of the visual coder. MaskVLM (Kwon et al., 2022) not only effectively integrates CLIP, MIM, and MLM, but also innovatively explores strategies for images and text to guide each other in generating masks, thus optimizing the synergy between modalities. However, these methods have limitations: they either don’t make full use of the available multimodal information, depending instead on features from a single mode to perform tasks, or their exploration of masking strategies is confined to using random methods.
2.3. Attention Masked modelling
@@
In the realm of attention-based mask modeling, hard patch mining (HPM) stands out as a groundbreaking training paradigm for MIM, designed to enhance a model’s comprehension of image content (Wang et al., 2023). HPM method first has the model generate a challenging mask and then trains the model to predict the masked patches, similar to traditional methods. This approach enables the model to learn where it is worth applying masks while simultaneously learning how to solve the problems. Also, HPM introduces new metrics to measure the difficulty of the pre-training task by reconstructing the loss, which could help the model to focus more on challenging problems.
Meanwhile, traditional random masking strategies, effective in general domains, fall short in medical contexts where precision is crucial (Chen et al., 2022; Huang et al., 2021; Müller et al., 2022; Zhang et al., 2022; Wu et al., 2023b; Zhou et al., 2022). MedIM (Xie et al., 2023) mitigates the unique challenges posed by healthcare data by improving mask strategies.
However, current methods either rely solely on single-mode features for their attention mechanisms (Kwon et al., 2022) or have limited interaction between multimodal features, such as only employing feature matching (Xie et al., 2023). Our proposed method solves these problems well by employing cross-attention for multimodal feature interaction while performing attention-masked modelling on features that incorporate multimodal information, which allows the model to understand image and text information in a more integrated way.
Figure 2. Our MMCLIP framework builds upon CLIP, integrating MIM and MLM modules, with a redesigned masking strategy to enhance model representation. The key contributions include three simple yet effective designs: generating masks through feature interaction, fusing masks to augment adaptability, and refining the text masking strategy with Medical Entity Recognition. Incorporating these designs into our multimodal pre-training framework significantly boosts the model’s zero-shot performance.
3. Methodology
Figure 2 illustrates the pipeline of the proposed MMCLIP. MMCLIP excels in representation learning by using image-report contrastive learning and masked modelling across both paired and unpaired medical image-report data. Central to MMCLIP are two innovative modules: Attention-Masked Image Modeling (AttMIM) and Entity-Driven Masked Language Modeling (EntMLM).
@@ 3. Methodology
EntMLM, on the other hand, concentrates on masking and reconstructing significant entities within medical reports, utilizing insights from image representations. This integration facilitates a more nuanced and detailed comprehension of medical terminology, enhancing the capabilities of both visual and textual learners within the MMCLIP framework.
3.1. Image and Text Encoders
3.1.1. Image encoder
We employ the Vision Transformer (ViT) (Dosovitskiy et al., 2020) with a patch size of 16 as the image encoder, initialized with the weights of the clip-base, to process the given input medical images $x$. Initially, $x$ is transformed into a sequence of flattened patches. These patches are subsequently embedded and introduced into a 12-layer transformer followed by a linear projection layer, denoted as $F(\cdot)$ and $Linear(\cdot)$, resulting in encoded representations of visual tokens:
(1) $E_{\text{img}}=Linear(F(x))\in\mathbb{R}^{N_{\text{img}}\times C},$
@@ 3.1.2. Text encoder
(2) $E_{\text{report}}=Linear(T(t_{report}))\in\mathbb{R}^{N_{\text{report}}\times C},$
(3) $E_{\text{prompt}}=Linear(T(t_{prompt}))\in\mathbb{R}^{N_{\text{prompt}}\times C},$
@@
Figure 3. The working mechanism of function M: The features obtained from different text data correspond to image features through the Cross Attention layer, querying the corresponding activated features to identify different regions of interest. These regions of interest are then combined, and a 25% subset is randomly selected to obtain the final Mask result. Unlike MediM, which only uses dot products to activate cross-modal response features, MMCLIP employs cross-attention for deeper feature activation. This allows the model to induce more precise key areas from text features.
3.2. Attention-masked Image Modeling Module
@@
3.2.1. Attentions extraction
@@
(4) $A_{r}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{report}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{report}},$
@@
(5) $A_{p}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{prompt}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{prompt}},$
@@
Image self-attention directs the model to focus on the most challenging parts of the image features during convergence, which tend to be more difficult regions (Wang et al., 2023), thus improving the model’s ability and efficiency in understanding the image (Zhang et al., 2024; Ji et al., 2024). Based on the self-attention mechanism, the image features after interacting with the image features themselves can be obtained as follows:
(6) $A_{i}=\text{softmax}\left(\frac{E_{\text{img}}E_{\text{img}}^{T}}{\sqrt{d_{k}}}\right)\cdot E_{\text{img}},$
@@
Figure 4. Different masking strategies result in varying concerns regarding mask application. Masks that are specifically tailored to the pathological characteristics of the lesion area are more effective than those applied randomly.
3.2.2. Attention-based mask generation and blending.
We introduce the function $M(A,\lambda)$, as shown in Fig. 3, which retrieves indices corresponding to the top values in feature matrix $A$, with the number of indices determined as a proportion $\lambda_{1}$ of the matrix’s size. This function effectively transforms these indexes into a mask corresponding to the input matrix, which is the high-activated attention masking results in the attention extraction. We blend the masks produced by different strategies in a union manner and define this result as $M_{b}$, formulated as:
(7) $M_{b}=M(Ar;\lambda_{1})\cup M(Ap;\lambda_{1})\cup M(Ai;\lambda_{1}).$
@@
(8) $\mathbb{M}_{\text{AttMIM}}={\rm RandomSubSet}(M_{b},\lambda_{2}).$
@@
(9) ${\rm MIM}_{\text{input}}=\mathbb{M}_{\text{AttMIM}}\cdot E_{\text{img}}.$
@@
3.3. Entity-driven Masked Language Modeling Module